# Annamaria Porreca Fabrizio Maturo Viviana Ventre

### v1

@@ 4 A brief introduction to Fuzzy Numbers
where $\mu_{A}(x)$ is a function, known as the membership function, that specifies the degree to which any element $x$ in $A$ belongs to the fuzzy set A. This implies a gradual transition from membership to non-membership [47].
Formally, a fuzzy set A is defined by its characteristic function:
@@
4.1 Definition of a fuzzy number
A fuzzy number $\mu(x)$ is a special case of a fuzzy set; in fact it can be defined as a fuzzy set, defined on real number, with a normal and convex membership function, such that there exists at least one point where the membership function takes the value “one”; it is a very useful tool for working with imprecise numerical quantities [45]. A fuzzy number is a function having as domain the set of real numbers and with values in $[0,1]$:
@@ 4.1 Definition of a fuzzy number
such that the following characteristics apply:
  1. 1.
    Bounded support: there are two real numbers a and b, with $a\leq b$, called the endpoints of $\mu$, such that:
    $\begin{cases}\mu(x)=0&\text{for}\qquad x\not\in[a,b]\\ \mu(x)>0&\text{for}\qquad x\in(a,b);\end{cases}$ (15)
  2. 2.
    Normality: there are two real numbers $c$ and $d$, with $a\leq c\leq d\leq b$ such that:
    $\mu(x)=1\qquad\text{if and only if}\qquad x\in[c,d].$ (16)
  3. 3.
    Convexity: $\mu(x)$ is a function increasing in the interval $[a,c]$ and decreasing in the interval $[d,b]$;
  4. 4.
    Compactness: for every $\alpha\in(0,1)$, the set $x\in\mathbb{R}:\mu(x)=\alpha$} is a closed interval.
The set of the real numbers $x$ such that $\mu(x)>0$ is said the support of the fuzzy number, and the interval $[c,d]$ is the core or central part. The intervals $[a,c)$ and $(d,b]$ are, respectively, the left part and the right part. The real numbers ${\mu}_{L}=c-a$, ${\mu}_{C}=d-c$, and ${\mu}_{R}=b-d$ are the left, middle, and right spreads, respectively. Their sum ${\mu}_{T}=b-a$ is the total spread of the fuzzy number [23]. The choice of the type of fuzzy numbers is delicate as it influences the characteristics of the processes of fuzzification and defuzzification of the outputs and the inputs [47]. The fuzzification allows to calculate the degree of membership of each numerical value assumed by one input variable to each fuzzy set defined for it. The defuzzification, instead, starting from the result obtained in the fuzzy inference process, calculates a real value for the output variable. The main membership functions representing fuzzy variables are triangular (TFN), trapezoidal, and bell-shaped. In this work, we concentrate on TFNs. A triangular fuzzy number $A$ is defined by the following membership function:
@@
Figure 1: Triangular fuzzy number.
4.2 Ranking fuzzy numbers
Given a universe set $X$, a fuzzy set $A_{i}$ on $X$ is defined as $\{(x,\mu_{A_{i}}(x))\mid x\in X\}$, where $\mu_{A_{i}}(x)$ denotes the membership degree of $x$ in $A_{i}$ [9]. The support of a fuzzy number is defined as:
$A_{i}=\operatorname{cl}\{x\in X\mid\mu_{A_{i}}(x)>0\},$
where $\operatorname{cl}$ represents the topological closure operator. Let $\mathcal{F}$ denote the set of all fuzzy numbers. Since a fuzzy number $A_{i}$ is fully characterized by its membership function values, denote this function by $A_{i}(x)$.
The primary ranking methods can be categorized into two types [9]. The first type maps fuzzy numbers directly onto the real line using transformation functions $M:\mathcal{F}\rightarrow\mathbb{R}$. These functions associate each fuzzy number with a real number and use the ordering $\geq$ on the real line. For most methods, a higher associated value indicates a higher rank:
$M(A_{i})\geq M(A_{j})\Rightarrow A_{i}\succeq_{M}A_{j},$
where $\succeq_{M}$ denotes the dominance relation induced by $M$.
The second type of ranking method generates fuzzy binary relations, defined as functions $M:\mathcal{F}\times\mathcal{F}\rightarrow[0,1]$, where $M(A_{i},A_{j})\in[0,1]$ measures the degree to which $A_{i}$ is greater than $A_{j}$. Accordingly, fuzzy numbers are ranked based on:
$M(A_{i},A_{j})\geq M(A_{j},A_{i})\Rightarrow A_{i}\succeq_{M}A_{j}.$
From the fuzzy binary relation among fuzzy numbers, certain procedures derive an ordering relation [38, 39]. One widely used ranking method is the Center of Gravity (CoG) of a fuzzy number, introduced by [30]:
$\operatorname{CoG}(A)=\frac{\int_{-\infty}^{\infty}xA(x)\,\mathrm{d}x}{\int_{-\infty}^{\infty}A(x)\,\mathrm{d}x}.$ (18)
There exist numerous other proposals in the literature for ranking fuzzy numbers [see, e.g., 42, 7, 16, 11, 12, 22, 1, 27]; however, ranking methods are secondary to our study focus and serve a functional role in our application. Hereafter, we employ the Center of Gravity, though our approach may extend to other strategies.
4.3 A brief introduction to Fuzzy Graphs (FGs)
Graphs are fundamental mathematical structures commonly used to model networks and relationships [5, 6]. Rosenfeld [34] introduced the concept of a fuzzy graph, applying fuzzy set theory to graphs. According to Rosenfeld [34], a fuzzy graph is defined as a graph composed of vertices and edges, where the edges have membership functions defined as follows:
Definition 4 (Rosenfeld [34])
A fuzzy graph $G=(\sigma,\mu)$ consists of functions $\sigma:S\rightarrow[0,1]$ and $\mu:S\times S\rightarrow[0,1]$, where for all $x,y\in S$, $\mu(x,y)\leq\min\{\sigma(x),\sigma(y)\}$.
This definition specifies that in a fuzzy graph, both vertices and edges have membership values in the interval $[0,1]$. Yeh and Bang [44] proposed an alternative formulation of fuzzy graphs tailored for clustering analysis:
Definition 5 (Yeh and Bang [44])
A fuzzy graph $G=(V,R)$ consists of a set of vertices $V$ and a fuzzy relation $R$ on $V$, where the edges between vertices have a membership function $\mu_{R}:V\times V\rightarrow[0,1]$.
Unlike Definition 4, Yeh and Bang’s formulation allows for both fuzzy vertices and fuzzy edges. They designed this version of fuzzy graphs to better suit clustering applications, providing a more flexible representation.
Since their inception, fuzzy graphs by Rosenfeld [34] and Yeh and Bang [44] have been extensively studied for various applications. Blue et al. [5, 6] categorized fuzzy graphs into different types, including:
@@ 4.3 A brief introduction to Fuzzy Graphs (FGs)
    Fuzzy sets of crisp graphs:
    $G=\{(G_{1},\mu(G_{1})),(G_{2},\mu(G_{2})),\ldots,(G_{n},\mu(G_{n}))\}$
    where each crisp graph $G_{i}$ is associated with a membership function $\mu(G_{i})$.
@@
    Crisp vertices and edges with fuzzy connectivity:
    $G=\{V,E\},\quad V=\{v_{1},v_{2},\ldots,v_{n}\},\quad E=\{(h_{i},\sigma_{i}),(t_{i},\tau_{i})\mid i=1,2,\ldots,m\}$
    The edges’ connectivity is fuzzy, where heads $h_{i}$ and tails $t_{i}$ have membership values in $[0,1]$, while vertices and edges themselves are crisp.
@@
    Fuzzy vertices and crisp edges:
    $G=\{V,E\},\quad V=\{(v_{1},\mu(v_{1})),(v_{2},\mu(v_{2})),\ldots,(v_{n},\mu(v_{n}))\},\quad E=\{e_{1},e_{2},\ldots,e_{n}\}$
@@
    The graph $G$ consists of fuzzy vertices with membership functions and crisp edges.
  4. 4.
    Crisp graph with fuzzy weight:
    $G=\{V,E\},\quad V=\{v_{1},v_{2},\ldots,v_{n}\},\quad E=\{e_{1},e_{2},\ldots,e_{n}\}$
    where edges have fuzzy weights defined by:
    $w_{i}=\{(w_{i,1},\mu(w_{i,1})),(w_{i,2},\mu(w_{i,2})),\ldots\}$
These classifications reflect the varying degrees of fuzziness applied to graphs, ranging from fully crisp vertices and edges to combinations of fuzzy vertices or edges with crisp counterparts.
4.4 Fuzzy operations
A fuzzy operation of $n$ variables is a function,
$f:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow[0,1]$
The most commonly used binary fuzzy operations are the “fuzzy intersections” or “t-norms” and the “fuzzy union” or “t-conorms”.
4.4.1 The “fuzzy intersection” or “t-norms”
Definition 6
Let $f:[0,1]\times[0,1]\rightarrow[0,1]$, it is a “fuzzy intersection” or “t-norm” if the following properties hold:
  • 1.
    (Commutative) $f(a,b)=f(b,a)$;
  • 2.
    (Associative) $f(f(a,b),c)=f(a,f(b,c))$;
  • 3.
    (Monotonicity) $a\leq a^{\prime}\Rightarrow f(a,b)\leq f(a^{\prime},b);\quad b\leq b^{\prime}\Rightarrow f(a,b)\leq f(a,b^{\prime})$;
  • 4.
    (Unity as a neutral element) $f(a,1)=a;\quad f(1,b)=b$.
We then write $afb$ to indicate $f(a,b)$. Some examples of fuzzy intersections are:
  • 1.
    Standard intersection $\cap$: $a\cap b=\min(a,b)$;
  • 2.
    Algebraic product $\odot$: $a\odot b=ab$;
  • 3.
    Limited difference $\odot_{d}$: $a\odot_{d}b=\max(0,a+b-1)$;
  • 4.
[…]
The following theorems hold:
Theorem 1
For each $a,b\in[0,1]$, $a\Delta b\leq a\odot b\leq a\cap b$.
Theorem 2
If $\otimes$ is any fuzzy intersection $\implies$ $a\Delta b\leq a\odot b\leq a\cap b$.
Theorem 3
The standard intersection is the only fuzzy intersection $\times$ which satisfies the following property:
  • 1.
    (Idempotence) $\forall a\in[0,1],a\odot a=a$.
4.4.2 The “fuzzy union” or “t-conorms”
Definition 7
Let $f:[0,1]\times[0,1]\rightarrow[0,1]$, it is a “fuzzy union” or “t-conorm” if the following properties hold:
  • 1.
    (Commutative) $f(a,b)=f(b,a)$;
  • 2.
    (Associative) $f(f(a,b),c)=f(a,f(b,c))$;
  • 3.
    (Monotonicity) $a\leq a^{\prime}\Rightarrow f(a,b)\leq f(a^{\prime},b);\quad b\leq b^{\prime}\Rightarrow f(a,b)\leq f(a,b^{\prime})$;
  • 4.
    (Zero as a neutral element) $f(a,0)=a;\quad f(0,b)=b$.
We then write $afb$ to indicate $f(a,b)$. Some examples of fuzzy unions are:
  • 1.
    Standard union $\cup$: $a\cup b=\max(a,b)$;
  • 2.
    Algebraic sum ssubscriptdirect-$\oplus_{s}$: $a\oplus_{s}b=a+b-ab$;
  • 3.
    Limited sum $\oplus_{L}$: $a\oplus_{L}b=\min(1,a+b)$;
  • 4.
[…]
The following theorems hold:
Theorem 4
For each $a,b\in[0,1]$, $a\cup b\leq a\oplus_{s}b\leq a\oplus_{L}b\leq a\Gamma b$.
Theorem 5
If $\oplus$ is any fuzzy union $\implies$ $a\cup b\leq a\oplus b\leq a\Gamma b$.
Theorem 6
The standard union is the only fuzzy union $\oplus$ which satisfies the following property:
  • 1.
    (Idempotence) $\forall a\in[0,1],a\oplus a=a$.
Definition 8
A fuzzy intersection $\otimes$ and a fuzzy union $\oplus$ are said to be associated if it holds: $a\otimes b+a\oplus b=a+b$ (Grassmann relation).
Theorem 7
The following fuzzy intersection and fuzzy union pairs are associated with each other:
  • 1.
    Standard Intersection and Standard Union;
  • 2.
    Algebraic product and Algebraic sum;
  • 3.
    Limited Difference and Limited Sum;
  • 4.
    Drastic intersection and Drastic union.
4.4.3 Fuzzy n-air operations
Let $h:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow[0,1]$ be an $n$-ary fuzzy operation. In order for it to represent an average of the values $v_{1},v_{2},\ldots,v_{n}$, it is necessary to require some axioms:
  • 1.
    (M1) (Idempotence)
    $(v_{1}=v_{2}=\ldots=v_{n}=k\in[0,1])\rightarrow h(v_{1},v_{2},\ldots,v_{n})=k$;
  • 2.
    (M2) (Growth)
    $(u_{i}\leq v_{i},i=1,2,\ldots,n)\rightarrow h(u_{1},u_{2},\ldots,u_{n})\leq h(v_{1},v_{2},\ldots,v_{n})$;
  • 3.
    (M3) (Commutativity)
    If $(u_{1},u_{2},\ldots,u_{n})$ is a permutation of $(v_{1},v_{2},\ldots,v_{n})$ then $h(u_{1},u_{2},\ldots,u_{n})=h(v_{1},v_{2},\ldots,v_{n})$;
  • 4.
    (M4) (Continuity)
    $h$ is a continuous function on $[0,1]^{n}$.
Definition 9
Let $w=(w_{1},w_{2},\ldots,w_{n})$ be a vector formed by elements belonging to the interval $[0,1]$ and such that $w_{1}+w_{2}+\ldots+w_{n}=1$. For each tuple $(a_{1},a_{2},\ldots,a_{n})$ of elements of $[0,1]$, let $(b_{1},b_{2},\ldots,b_{n})$ be the permutation of $(a_{1},a_{2},\ldots,a_{n})$ satisfying the condition $b_{1}\geq b_{2}\geq\ldots\geq b_{n}$. The ordered weighted average of $a_{1},a_{2},\ldots,a_{n}$, weighted with the weight vector $w$ (OWA = Ordered Weighted Average), is the number:
$m_{w}(a_{1},a_{2},\ldots,a_{n})=b_{1}w_{1}+b_{2}w_{2}+\ldots+b_{n}w_{n}$
The following theorems hold:
Theorem 8
$m_{w}(a_{1},a_{2},\ldots,a_{n})\in[0,1]$.
Theorem 9
For any weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ that is non-negative and such that $w_{1}+w_{2}+\ldots+w_{n}=1$, the operation $h:(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}\rightarrow m_{w}(a_{1},a_{2},\ldots,a_{n})\in[0,1]$ satisfies properties (M1), (M2), (M3), and (M4).
Theorem 10
Let $w=(w_{1},w_{2},\ldots,w_{n})$ and $w^{\prime}=(w_{1},w_{2},\ldots,w_{i}+h,\ldots,w_{j}-h,\ldots,w_{n})$ with $h>0$, $i<j$, two non-negative weight vectors such that $w_{1}+w_{2}+\ldots+w_{n}=1$. Then for all $(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}$, $m_{w}(a_{1},a_{2},\ldots,a_{n})\leq m_{w^{\prime}}(a_{1},a_{2},\ldots,a_{n})$.
Theorem 11
Let $w_{\text{max}}$ be the weight vector with the first component equal to 1 and all others zero, and let $w_{\text{min}}$ be the weight vector with the last component equal to 1 and all others zero. For any weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ that is non-negative and such that $w_{1}+w_{2}+\ldots+w_{n}=1$, it holds that for all $(a_{1},a_{2},\ldots,a_{n})\in[0,1]^{n}$,
$m_{w_{\text{min}}}(a_{1},a_{2},\ldots,a_{n})\leq m_{w}(a_{1},a_{2},\ldots,a_{n})\leq m_{w_{\text{max}}}(a_{1},a_{2},\ldots,a_{n}).$
Definition 10
Let $\diamondsuit$ be a binary operation on $[0,1]$ which satisfies the commutative, associative, and monotonic properties. The $n$-ary operation generated by $\diamondsuit$ is the function:
$\diamondsuit_{n}:(v_{1},v_{2},\ldots,v_{n})\in[0,1]^{n}\rightarrow v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}\in[0,1].$
We observe that based on the commutative and associative properties, $v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}$ does not depend on the order in which the operations are carried out or on any permutations of the elements. Additionally, properties (M2) and (M3) apply. If $\diamondsuit$ is also idempotent (M1) and continuous (M4), then the $n$-ary operation generated by $\diamondsuit$ is a weighted ordered average (OWA). This occurs if there exists a weight vector $w=(w_{1},w_{2},\ldots,w_{n})$ such that:
$v_{1}\diamondsuit v_{2}\diamondsuit\ldots\diamondsuit v_{n}=m_{w}(v_{1},v_{2},\ldots,v_{n}).$
Theorem 12
If $\diamondsuit$ is the standard intersection $\cap$, then $\diamondsuit_{n}=v_{1}\cap v_{2}\cap\ldots\cap v_{n}$ is a weighted ordered average associated with the weight $w_{\text{min}}=(0,0,\ldots,1)$, i.e., $m_{w_{\text{min}}}(v_{1},v_{2},\ldots,v_{n})$.
Theorem 13
If $\diamondsuit$ is the standard union $\cup$, then $\diamondsuit_{n}=v_{1}\cup v_{2}\cup\ldots\cup v_{n}$ is a weighted ordered average associated with the weight $w_{\text{max}}=(1,0,\ldots,0)$, i.e., $m_{w_{\text{max}}}(v_{1},v_{2},\ldots,v_{n})$.
An example of a weighted ordered average not generated by a binary operation is the simple arithmetic average where $m_{w}(v_{1},v_{2},\ldots,v_{n})=\frac{v_{1}+v_{2}+\ldots+v_{n}}{n}$, and therefore $w=\left(\frac{1}{n},\frac{1}{n},\ldots,\frac{1}{n}\right)$.
In fact, if $\diamondsuit$ is the binary operation that associates the mean $\frac{a+b}{2}$ to each ordered pair $(a,b)$ of elements of $[0,1]$, then the commutative, monotonic, continuous, and idempotent properties hold, but the associative property does not.
4.5 Ordering and metric space structures of the OWA operators
4.5.1 Sorting structure
Let $W$ be the set of vectors $w=(w_{1},w_{2},\ldots,w_{n})$ formed by $n$ elements belonging to the interval $[0,1]$ such that $w_{1}+w_{2}+\ldots+w_{n}=1$. In $W$, an order $\leq$ can be defined for each pair $w=(w_{1},w_{2},\ldots,w_{n}),w^{\prime}=(w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{n})$ as follows:
$w\leq w^{\prime}\Leftrightarrow(\forall i\in\{1,2,\ldots,n-1\},w_{i}\leq w^{\prime}_{i})$ (19)
Theorem 14
The pair $(W,\leq)$ forms a lattice with at least $w_{\min}=(0,0,\ldots,1)$ and maximum $w_{\max}=(1,0,\ldots,0)$.
4.5.2 Structure of the metric space
Let $W$ be the set of vectors $w=(w_{1},w_{2},\ldots,w_{n})$ formed by $n$ elements belonging to the interval $[0,1]$ such that $w_{1}+w_{2}+\ldots+w_{n}=1$. Let $k=(k_{1},k_{2},\ldots,k_{n})$ be a vector satisfying $k_{1}>k_{2}>\ldots>k_{n}$, with $k_{1}=1$ and $k_{n}=0$. In $W$, a distance (or metric) $d$, associated with $k$, can be defined for each $w,w^{\prime}\in W$ as:
$d(w,w^{\prime})=k_{1}|w_{1}-w^{\prime}_{1}|+k_{2}|w_{2}-w^{\prime}_{2}|+\ldots+k_{n}|w_{n}-w^{\prime}_{n}|$ (20)
Some authors [43, 8] have proposed the expression
$k_{i}=\frac{n_{i}}{n-1}$ (21)
Thus, (20) can be rewritten as:
$d(w,w^{\prime})=\frac{\left[\sum_{i=1}^{n-1}(n-i)|w_{i}-w^{\prime}_{i}|\right]}{n-1}$ (22)
Theorem 15
Let $d$ be the distance defined by (20). The pair $(W,d)$ forms a metric space. Furthermore,
  • 1.
    (D1) $d(w,w^{\prime})\leq 1$;
  • 2.
    (D2) $d(w,w^{\prime})=1\iff\{w,w^{\prime}\}=\{w_{\min},w_{\max}\}$;
  • 3.
    (D3) For all $w\in W$, $d(w,w_{\min})+d(w,w_{\max})=1$.
According to Theorem 15, we can define a ”degree of proximity” $v(w,w^{*})$ of a vector $w$ from a fixed vector $w^{*}$ by setting $v(w,w^{*})=1-d(w,w^{*})$. In particular, we can consider:
  • 1.
    The proximity $v(w,w_{\max})$ of a vector $w=(w_{1},w_{2},\ldots,w_{n})$ from the vector $w_{\max}$, called orness (unionity). If we assume (21), it results in:
    $\text{orness}(w)=v(w,w_{\max})=\left[\sum_{i=1}^{n-1}\frac{(n-i)w_{i}}{n-1}\right]$ (23)
    With simple calculations, we find:
    $\text{orness}(w_{\max})=1,\quad\text{orness}(w_{\min})=0,\quad\text{orness}(w_{\text{med}})=0.5$
    where $w_{\text{med}}$ is the vector having all components equal to $1/n$.
  • 2.
    The proximity $v(w,w_{\min})$ of a vector $w=(w_{1},w_{2},\ldots,w_{n})$ from the vector $w_{\min}$, called andness (intersection). From (D3), it follows that:
    $\text{andness}(w)=1-\text{orness}(w)$ (24)
    If we assume (21), then it results in:
    $\text{andness}(w)=1-\left[\sum_{i=1}^{n-1}\frac{(n-i)w_{i}}{n-1}\right]$ (25)
  • 3.
    The mediality of a vector of weights $w$, defined as $v(w,w_{\text{med}})$, where $w_{\text{med}}$ is the vector having all components equal to $1/n$.
4.5.3 FOWA: The Fuzzy Ordered Weighted Averaging Operator
Recently, Chen and Chen [13] formulated the FOWA operator to deal with information expressed by fuzzy numbers.
Definition 11
[37] The fuzzy ordered weighted averaging operator, denoted by FOWA, is the mapping $\text{FOWA}:\widetilde{\mathscr{P}}(\mathbb{R}^{+})^{n}\rightarrow\widetilde{\mathscr{P}}(\mathbb{R}^{+})$, which has an associated weighting vector $(w_{1},\ldots,w_{n})$, such that
$\text{FOWA}(\tilde{x}_{1},\ldots,\tilde{x}_{n})=\bigoplus_{i\in[n]}(w_{i}\odot\tilde{x}_{(i)})$
where the operations $\oplus$ and $\odot$ are a t-conorm and a t-norm respectively, and ()($\cdot$)( ⋅ ) denotes a permutation on $[n]$ such that $\tilde{x}_{(1)}\leq\ldots\leq\tilde{x}_{(n)}$.
It is evident that the FOWA operator extends the OWA operator. Particularly, for a weight vector $w$, we have:
  • 1.
    If $\mathbf{w}=(1,0,\ldots,0)$, then FOWA $=\tilde{x}_{(1)}$ (fuzzy minimum).
  • 2.
    If $\mathbf{w}=(0,\ldots,0,1)$, then FOWA $=\tilde{x}_{(n)}$ (fuzzy maximum).
  • 3.
    If $w_{i}=\frac{1}{n}$ for all $i$, then FOWA $=$ FAA (fuzzy arithmetic averaging).
If FOWA concerns triangular fuzzy numbers, it is simply referred to as FTOWA.
5 Fuzzy Graphs: Generalization of Weighted Graphs
In this work, we propose the use of Fuzzy Graphs (FG) as a generalization of Weighted Graphs (WG). A WG is one in which edges are associated with weights. A WG can be represented as $G(V,E,W)$, where $W$ represents the weights associated with each edge ($|W|=|E|$). For an adjacency matrix representation, instead of 00 and $1$, we can use the weight associated with the edge. This saves space by combining $E$ and $W$ into one adjacency matrix $A$, assuming an edge exists between $v_{i}$ and $v_{j}$ if and only if $W_{i,j}\neq 0$.
In the literature, there are different definitions and proposals for the concept of fuzzy graph [see e.g. 49]. Here, we focus on graphs where only the edges are fuzzy. Thus, FGs are considered as a generalization of WG, where edges are associated with fuzzy numbers.
@@ 5 Fuzzy Graphs: Generalization of Weighted Graphs
Assume that $v_{0}A_{1}v_{1}A_{2}v_{2}\ldots A_{k}v_{k}$ is a path from $v_{0}$ to $v_{k}$ in $\widetilde{G}_{dir}$, then $\widetilde{s}_{d}(\widetilde{\omega})=\bigwedge_{i=1}^{k}\mu\left(A_{i}\right)$ is called the directed fuzzy intensity of path $\widetilde{\omega}$.
In Definition 2.3, $\mu(e_{i})$ is a membership function.
@@
If $\bar{\omega}_{k}$ are $n$ paths from $u$ to $v$ in $\widetilde{G}_{dir}$, then $\widetilde{s}_{d}(u,v)=\bigvee_{k=1}^{n}\widetilde{s}_{d}(\widetilde{\omega}_{k})$ is called the directed fuzzy connected intensity from $u$ to $v$.
Here, $k=1,2,\ldots,n$. If there is no path from $u$ to $v$, then $\widetilde{s}_{d}(u,v)=0$. If $u=v$, then $\widetilde{s}_{d}(u,v)=1$. In UFSN, $\bar{s}(u,v)$ always equals $\widetilde{s}(v,u)$. However, $\widetilde{s}_{d}(u,v)$ is not always equal to $\widetilde{s}_{d}(v,u)$ in DFSN.

### v2

@@ 3.1 Fuzzy Numbers
where $\mu_{A}(x)$ is a function, known as the membership function, that specifies the degree to which any element $x$ in $A$ belongs to the fuzzy set A. This implies a gradual transition from membership to non-membership [Zadeh, 1975]. Formally, a fuzzy set A is defined by its characteristic function:
@@
A fuzzy number $\mu(x)$ is a special case of a fuzzy set; in fact it can be defined as a fuzzy set, defined on real number, with a normal and convex membership function, such that there exists at least one point where the membership function takes the value “one”; it is a very useful tool for working with imprecise numerical quantities [Zadeh, 1965, Maturo et al., 2020, Maturo and Maturo, 2017, Maturo, 2016]. A fuzzy number is a function having as domain the set of real numbers and with values in $[0,1]$:
@@
such that the following characteristics apply. 1) Bounded support: there are two real numbers a and b, with $a\leq b$, called the endpoints of $\mu$, such that $\mu(x)=0$ for $x\not\in[a,b]$ and $\mu(x)>0$ for $x\in(a,b)$; 2) Normality: there are two real numbers $c$ and $d$, with $a\leq c\leq d\leq b$ such that $\mu(x)=1$ if and only if $x\in[c,d]$; 3) Convexity: $\mu(x)$ is a function increasing in the interval $[a,c]$ and decreasing in the interval $[d,b]$; 4) Compactness: for every $\alpha\in(0,1)$, the set $x\in\mathbb{R}:\mu(x)=\alpha$} is a closed interval. The set of the real numbers $x$ such that $\mu(x)>0$ is said the support of the fuzzy number, and the interval $[c,d]$ is the core or central part. The intervals $[a,c)$ and $(d,b]$ are, respectively, the left part and the right part. The real numbers ${\mu}_{L}=c-a$, ${\mu}_{C}=d-c$, and ${\mu}_{R}=b-d$ are the left, middle, and right spreads, respectively. Their sum ${\mu}_{T}=b-a$ is the total spread of the fuzzy number [Klir, 2006].
The choice of the type of fuzzy numbers influences the characteristics of the processes of fuzzification and defuzzification of the outputs and the inputs [Zadeh, 1975]. The primary membership functions used to represent fuzzy variables include triangular (TFN), trapezoidal, and bell-shaped functions. In this work, we focus specifically on TFNs. A TFN $A$ is defined by the following membership function:
@@
Figure 1: Triangular fuzzy number (TFN).
3.2 Basic Pseudo-Operations on Triangular Fuzzy Numbers
In the context of fuzzy logic, it is often necessary to extend classical arithmetic operations to fuzzy numbers. These extended operations are referred to as pseudo-operations. Specifically, pseudo-operations allow us to perform arithmetic in a fuzzy environment[Simo and Gwét, 2018]. Given two triangular fuzzy numbers $\tilde{A}=(a_{l},a_{m},a_{r})$ and $\tilde{B}=(b_{l},b_{m},b_{r})$, the following pseudo-operations are defined:
@@ 3.2 Basic Pseudo-Operations on Triangular Fuzzy Numbers
    Addition ($\oplus$):
    $\tilde{A}\oplus\tilde{B}=(a_{l}+b_{l},a_{m}+b_{m},a_{r}+b_{r})$ (5)
@@
    Multiplication ($\otimes$):
    $\tilde{A}\otimes\tilde{B}=(a_{l}\cdot b_{l},a_{m}\cdot b_{m},a_{r}\cdot b_{r})$ (6)
@@
    Scalar Multiplication ($\odot$):
    $w\odot\tilde{A}=(w\cdot a_{l},w\cdot a_{m},w\cdot a_{r})$ (7)
3.3 Fuzzy Triangular Ordered Weighted Arithmetic (FTOWA)
The Fuzzy Triangular Ordered Weighted Arithmetic (FTOWA) Operator is an aggregation operator specifically designed for triangular fuzzy numbers. The FTOWA operator aggregates a set of triangular fuzzy numbers by considering their ordered weights in a fuzzy environment[Simo and Gwét, 2018]. It is defined as:
$\text{FTOWA}(\tilde{A}_{1},\tilde{A}_{2},\dots,\tilde{A}_{n})=\bigoplus_{i=1}^{n}w_{i}\odot\tilde{A}_{(i)}$ (8)
where $\tilde{A}_{(i)}=(a_{(i)l},a_{(i)m},a_{(i)r})$ is the $i$-th largest TFN in the ordered set, and $w_{i}$ is the weight associated with the $i$-th ordered fuzzy number.
The FTOWA operator includes various specific cases depending on the choice of weights $w_{i}$:
  1. 1.
    Fuzzy Minimum: If $w=(1,0,\dots,0)$, then FTOWA gives $\tilde{A}_{(1)}$, the smallest triangular fuzzy number.
  2. 2.
    Fuzzy Maximum: If $w=(0,\dots,0,1)$, then FTOWA gives $\tilde{A}_{(n)}$, the largest triangular fuzzy number.
  3. 3.
    Fuzzy Arithmetic Average: If $w_{i}=\frac{1}{n}$ for all $i$, then FTOWA reduces to the arithmetic average of the triangular fuzzy numbers.
The FTOWA operator preserves essential properties like monotonicity, continuity, and idempotency, making it a versatile tool in decision-making processes where the information is represented as triangular fuzzy numbers.
3.4 Fuzzy Graphs
Rosenfeld [1975] introduced fuzzy graphs by applying fuzzy set theory, where both vertices and edges are characterized by membership values in $[0,1]$. Formally:
Definition 1 (Rosenfeld [Rosenfeld, 1975])
A fuzzy graph $G=(\sigma,\mu)$ consists of functions $\sigma:S\rightarrow[0,1]$ and $\mu:S\times S\rightarrow[0,1]$, where for all $x,y\in S$, $\mu(x,y)\leq\min{\sigma(x),\sigma(y)}$.
Yeh and Bang [1975] later extended this concept for clustering analysis, allowing both vertices and edges to be fuzzy:
Definition 2 (Yeh and Bang [Yeh and Bang, 1975])
@@ 3.4 Fuzzy Graphs
A fuzzy graph $G=(V,R)$ consists of a set of vertices $V$ and a fuzzy relation $R$ on $V$, where edges have a membership function $\mu_{R}:V\times V\rightarrow[0,1]$.
Subsequently, Blue et al. [1997, 2002] categorized fuzzy graphs into types, such as fuzzy sets of crisp graphs, crisp vertices and edges with fuzzy connectivity, fuzzy vertices with crisp edges, and crisp graphs with fuzzy weights. These variations reflect different levels of fuzziness in graph structures, from fully crisp to various combinations of fuzzy elements.
In this paper, we focus specifically on the case where the ties (or edges) in the graph are fuzzy, while the actors (or vertices) are crisp. This approach allows us to model the uncertainty or varying strength of relationships between clearly defined actors, reflecting more realistic scenarios in social network analysis where connections may not be strictly binary but can vary in intensity or reliability. We propose the use of Fuzzy Graphs (FG) as a generalization of Weighted Graphs (WG). A WG is one in which edges are associated with weights. A WG can be represented as $G(V,E,W)$, where $W$ represents the weights associated with each edge ($|W|=|E|$). For an adjacency matrix representation, instead of 00 and $1$, we can use the weight associated with the edge. This saves space by combining $E$ and $W$ into one adjacency matrix $A$, assuming an edge exists between $v_{i}$ and $v_{j}$ if and only if $W_{i,j}\neq 0$. Thus, FGs are considered as a generalization of WG, where edges are associated with fuzzy numbers.
@@
Suppose $v_{0}A_{1}v_{1}A_{2}v_{2}\dots A_{k}v_{k}$ is a path from node $v_{0}$ to $v_{k}$ in a directed fuzzy social network (denoted as $\widetilde{G}_{\text{dir}}$). The directed fuzzy intensity of this path, denoted as $\tilde{s}_{d}(\tilde{\omega})$, is calculated as the minimum membership function value among all the edges in the path:
$\tilde{s}_{d}(\tilde{\omega})=\bigwedge_{i=1}^{k}\mu(A_{i})$
where $\mu(e_{i})$ represents the membership function associated with the edge $e_{i}$ in a fuzzy graph or network.
@@
If there are $n$ paths from node $u$ to node $v$ in the directed fuzzy social network $\widetilde{G}_{\text{dir}}$, the directed fuzzy connected intensity from $u$ to $v$ is defined as:
$\widetilde{s}_{d}(u,v)=\bigvee_{k=1}^{n}\widetilde{s}_{d}(\widetilde{\omega}_{k})$
where $\widetilde{\omega}_{k}$ represents each path from $u$ to $v$, and $\bigvee$ indicates the maximum (supremum) of these intensities.
If there is no path from $u$ to $v$, $\widetilde{s}_{d}(u,v)=0$. If $u=v$, $\widetilde{s}_{d}(u,v)=1$. In UFSN, the connected intensity from $u$ to $v$ is equal to the intensity from $v$ to $u$. However, in DFSN, $\widetilde{s}_{d}(u,v)$ is not necessarily equal to $\widetilde{s}_{d}(v,u)$.